OpenStudy (savannah_noelle):

Algebra 2, 5 multiple choice questions, help please :(

OpenStudy (savannah_noelle):

1. Given the function f(x)=x^3+x^2-2x+1, what is resulting function when f(X) is shifted to the left 1 unit? A. f(x-1)=x^3-2x^2-x+3 B. f(x+1)=x^3+4x^2+3x+1 C. f(x)-1=x^3+x^2-2x D. f(x)+1=x^3+x^2-2x+2 2. Calculate the average rate of change for function f(x)=x^4+3x^3-5x^2+2x-2, from x=-1 to x=1. A.-1 B. 1 C.-5 D. 5 3.Without using technology, describe the end behavior of f(x)=-3x^38+7x^3-12x+13. A. Down on the left, down on the right. B. Down on the left, up on the right. C. Up on the left, down on the right. D. Up on the left, up on the right. 4.Given the function f(x)=-3x^3+9x^2-2x+3, what part of the function indicates that the left end starts at the top of the graph? A. The degree of the first term B. The degree of the last term C. The coefficient on the first term D. The coefficient of the last term 5. Graph the function f(x)=-x^4+7x^3-9x^2-3x+10 using graphing technology and identify for which values of x the graph is decreasing. A. From x=-2 to x=-1 B. From x=0 to x=1 C. From x=1 to x=2 D. x=2 to x=4

OpenStudy (aggie12341):

A, B, D, A, C

OpenStudy (savannah_noelle):

thank you!

OpenStudy (aggie12341):

medal!

OpenStudy (savannah_noelle):

wow...none of those answers were correct and I got a 0%..so no medal.

OpenStudy (savannah_noelle):

Not even one of those answers was correct @aggie12341

OpenStudy (aggie12341):

Yea it was C, D B A D

OpenStudy (savannah_noelle):

I feel like you're just making up answers, lol not cool.

OpenStudy (aggie12341):

Approximately two summers ago I was accepted to the Summer Pre-Graduate Research Experience (SPG RE) program offered by the Graduate School at the University of North Carolina. The program usually gives preference to rising seniors, but an exception was made in my case allowing me to attend as a junior. I was matched to a mentor, XXXXX XXXXX PhD, in the Department of Biostatistics. Professor XXXXX, knowing my limited background, introduced me to biostatistics by letting me observe his activities as a biostatistician and teaching me basic statistical analyses and techniques. Professor XXXXX introduced me to these concepts from a calculus point of view to make it applicable and easy to understand. From that initial exposure I became excited to find a field that allowed me to combine my mathematical background to the medical sciences. I was also excited to find the diverse range of research areas to which biostatisticians can apply their skill. Consequently, I was invited to participate in the program the following summer with Professor XXXXX as my mentor again. My experiences each summer were vastly different, and once again, I was attracted to the diversity of the field. The first summer, I was very inexperienced, so my responsibilities included learning basic statistical techniques, computing skills (SAS, word processing, etc.), communication skills, and completing a research project involving measurement error variability of pulmonary lung functions. However, during the second summer, I was given more responsibilities. I was one of the acting statistical consultants to a project involving violence among African American adolescents that was conducted by another student, YYYYY in SPGRE. YYYYY's mentor was Professor ZZZZZ in the Department of Maternal and Child Health at UNC. In addition to completing our project this summer, we were responsible for teaching two new students SAS programming skills, data management, performing preliminary analysis on the larger umbrella study from which ours was drawn, and managing the progress for our own project. Even though I enjoyed both summer experiences, I found my second summer to be more invigorating than the first.